OpenStudy (anonymous):

y = 2x – 1 and y = x + 3 (1 point) (4, 7) (7, 4) (–7, –4) infinite solutions

OpenStudy (anonymous):

2x-1= x+3 2x-x= 4 You should be able to solve now, right?

OpenStudy (ryan123345):